Aimee mows lawns at a rate of 30 square feet per minute, Judy mows at 40 square feet per minute, and sue mows 60 square feet per minute. They accept a job that requires they mow 22,400 square feet of lawn. Aimee begins mowing at 7 am but Judy does not begin for another 15 minutes. Sue begins mowing at 8:45am and they finish together. To the nearest minute, how long did Judy work?
Answer by edjones(7569)

By 8:45 Aimee mowed 30(60+45)=30*105=3150 ft^2
By 8:45 Judy mowed 40(60+30)=30*90=2700 ft^2
22400-3150-2700=16550 left to mow
30+40+60=130/min combined
16550/130=127 min left to mow.
127+90=217 min. Judy worked.
.
Ed